# Warehouse storage and retrieval optimization via clustering, dynamic systems modeling, and GPU-accelerated routing

### v1

@@ 1. Introduction
The system employs clustering on two levels: (1) Order clustering, which reorganizes storage layouts to concentrate frequently co-ordered items and minimize long-term travel distance, and (2) Picking node clustering, which reduces computational complexity in route optimization while maintaining near-optimal results. These techniques are analyzed mathematically in Sections 2 and 3 and evaluated through numerical experiments in Section 4. By continuously adapting storage configurations and applying combinatorial optimization strategies, this approach balances short-term efficiency gains with long-term structural improvements, ensuring that warehouse operations remain effective even amid fluctuating order patterns.
@@ 2.1. The state space
We consider a situation where $N$ types of articles $a_{1},a_{2},\ldots,a_{N}$ are stored in a warehouse at nodes arranged in three-dimensional space representing racks with shelves at different heights. Each shelf on a rack constitutes its own node, and each node is identified with an element in a three-dimensional array having integer coordinates
@@
where =$\mathbb{N}=\{0,1,2,\ldots\}$ is the set of natural number, and $n_{x}\times n_{y}\times n_{z}$ is the total number of nodes in the warehouse. We assume there is at most one article type $a_{n}$ at each node (some nodes being empty), and identify each article type $a_{n}$ with the corresponding ordinal $n\in[1,N]$. Empty nodes are assigned article type 00.
For each article type $n\in[1,N]$, let $M_{n}$ be the maximum number of parcels that can be stored at an individual node. We assume this number is independent of which node is being considered; it only depends on what type of article is stored at that node. Introduce the space $\mathcal{A}=[0,N]^{n_{x}\times n_{y}\times n_{z}}$ of three-dimensional arrays. Each array $A=(a_{ijk})_{ijk}\in\mathcal{A}$ corresponds to one particular choice of how to distribute article types in the warehouse. (The array with only zeros represents an empty warehouse with no article types, and the array with only ones is a warehouse that only contains articles of type $a_{1}$.) Each array element $a_{ijk}$ is thus an integer $n$ between 00 and $N$, representing the fact that the node at $(i,j,k)$ contains article type $a_{n}$.
@@ 2.2. Orders
where $m_{n}\in\mathbb{N}$ is the number of parcels of article $a_{n}$ contained in the order. (Here we shall by an order mean any order – individual customer orders or a collection of customer orders – that has been decided should be picked from the warehouse.) Given a state $x=(A,M)$ and an order $o$, the Order Processing algorithm (defined in Algorithm 1 in Appendix A below) and a route optimization algorithm (which involves the proposed clustering step based on similarity and proximity, see §2.4) provide the nodes from which the order should be collected, along with the picking route between these nodes. The first part is described by a map $w$ given by
@@
The Check Stock algorithm (Algorithm 3) verifies if there is sufficient stock to fill the order without picking the last parcel on a shelf, i.e., whether $m_{ijk}>w_{ijk}$ for all $(i,j,k)$. If this is not the case, i.e., if $m_{ijk}\leq w_{ijk}$ for some $(i,j,k)$ such that $w_{ijk}\geq 1$, then the Check Stock and Move Articles algorithm of the warehouse optimization routine (Algorithm 4) computes a new warehouse location $(i^{\prime},j^{\prime},k^{\prime})$ at which to restock article $a_{ijk}$ (collecting from the new location after emptying the balance at the old location), and the article array $A$ is updated to $A^{\prime}$ in which $a_{ijk}$ is replaced by 00, and $a_{i^{\prime}j^{\prime}k^{\prime}}$ is changed from 00 to $a_{ijk}$. This gives an updated state
@@ 2.3. Random dynamical systems
We now recall the notion of a random dynamical system (RDS), a symbiosis of ergodic theory and topological dynamical systems. An advantage of the RDS framework is that it also applies in situations where orders have a random element. Standard references are Arnold [2] and Strogatz [17].
We use discrete time $n\in\mathbb{N}$, where =$\mathbb{N}=\{0,1,2,\ldots\}$ as before is the set of natural numbers. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We take $\Omega$ to be the shift space consisting of sequences $\omega=(o_{1},o_{2},\ldots)$ of orders $o_{n}$, and let $\theta:\Omega\to\Omega$ to be the shift operator $\theta\omega=(o_{2},o_{3},\ldots)$. This models a sequence of orders $\omega\in\Omega$, placed with probability $\mathbb{P}$, coming in to the warehouse to be filled one after the other, with $\theta$ driving the flow of orders. Let $X$ be the state space describing different states of the warehouse. Let $F:\omega\mapsto F_{\omega}$ be a random variable from $\Omega$ to the set of maps on $X$, i.e., for all $\omega\in\Omega$, $F_{\omega}$ is a map $F_{\omega}:X\to X$. Consider the skew product defined by $\Theta(\omega,x)=(\theta\omega,F_{\omega}(x))$. Iterating gives
@@ 2.4. Routes and clusters
At stopping nodes with more than one picking node, we can pre-determine the order in which picking is carried out – e.g., top to bottom – so this does not have to not included in the route optimization. We will therefore not distinguish between picking routes and routes between stopping nodes.
Figure 1. Each of the top panels show one of three clusters resulting from $K$-means clustering applied to an order with 200 article types given a randomized initial warehouse state $x_{0}$. The nodes in the warehouse have coordinates $(i,j,k)$ with $i,j,k=1,\ldots,10$. Stopping nodes are indicated with color coding describing how many picking nodes a given stopping node contains. Each of the bottom panels show one of three clusters resulting from $K$-means clustering applied to the same order being picked for the 100:th time, that is, given the warehouse state $x_{99}$ as in (5), obtained after 99 picking iterations. In the top panels, all three clusters are spread out over the entire warehouse with much overlap between clusters. In contrast, the clusters in the bottom panels are concentrated and much better separated.
@@
Remark.
Note that we define clusters mathematically as a subset of the picking nodes for a given order, as opposed to dividing the picking nodes from a collection of orders into different groups or clusters. But since an order here can refer to the total collection of individual customer orders to be picked at a certain time our definition covers also the second scenario.
@@ 3. Route Optimization
Suppose first that an undirected route with two boundary nodes has been found for each cluster. We determine the computational complexity of finding the best global route going through all clusters along these routes using a combinatorial argument. First, a starting cluster is chosen which can be done in $m$ ways. From that cluster, a global starting point should be determined from the pair of that cluster’s boundary nodes, which can be done in $2$ ways. Next, a second cluster is chosen which can be done in $m-1$ ways. From that cluster, a starting point connecting with the first cluster’s end point should be determined from the second cluster’s pair of boundary nodes, which can be done in $2$ ways. Continuing this way we see that there are $m!\times 2^{m}$ ways to choose a global route with a start and end. Hence, there are half as many global undirected routes, namely $m!\times 2^{m-1}$.
@@
Figure 2. An order with 15 stopping positions clustered into 3 regions (left), and the reduced routes for each cluster represented by two boundary nodes (right).
@@
Of course, we have been comparing the results of the theorem above to a brute force approach for finding optimal routes without employing sophisticated algorithms to reduce computational complexity. However, using for example the Held & Karp algorithm [11] reduces the time complexity from $\mathcal{O}(n!)$ to $\mathcal{O}(n^{2}\cdot 2^{n})$, which, with $n=12$, is still 3 orders of magnitude higher than the examples above – and this without using a corresponding algorithm to find optimal routes within clusters. Hence, the gains in time complexity from the approximate method are still substantial.
@@ 4. Numerical experiments
    The number of article types is $N=890$, and we assume that the maximum number of articles of each article type that a node can stock is $10$, so $M_{a_{ijk}}=10$ for all $(i,j,k)$. Thus, the state space $X$ is
@@
    The initial warehouse state $x_{0}$ is randomized with the following constraints: It contains $11$ $xy$-coordinate pairs $(i_{0},j_{0}),(i_{1},j_{1}),\ldots,(i_{10},j_{10})$ such that the nodes $(i_{0},j_{0},k),\ldots(i_{10},j_{10},k)$ are empty for all $k\in[1,10]$. This corresponds to $11$ empty racks, for a total of $110$ empty nodes. A unique article type $a_{ijk}\in[1,890]$ is assigned to each of the remaining $890$ nodes, and we assume that the initial warehouse state is fully stocked, so the balance at each nonempty node $10$.
@@
    An order $o$ to be picked is a collection of $20$ individual customer or purchase orders. A purchase order consists of $10$ article types with between $1$ and $10$ parcels of each type. The $20$ purchase orders are mutually disjoint, so $o$ consists of $200$ article types. In each picking iteration, the $20$ purchase orders are divided into three groups using $K$-means clustering based on where the articles are stored in the warehouse at that time. Each such group is a cluster. An example can be seen in Figure 1.
@@ 4.1. Experiment 1
Figure 3. Three snapshots of cluster distributions from experiment 1, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n$ (i.e., in the language of §2.4 they show the clusters computed for the $n$:th order), with $n=1$ (left), $n=10$ (middle), $n=100$ (right). The orders are the same in every iteration. The figure clearly shows that as time progresses, the separation between cluster centers increases, while the bumps become more articulated due to a decreased variance. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_unperturbed/0_3r1qr1w4.
@@ 4.2. Experiment 2
Figure 4. Three snapshots of cluster distributions from experiment 2, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $10\%$ on average. Although not as pronounced, we still clearly see an increased separation and decreased variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_static_perturbation/0_t9n5lbme.

### v2

# Warehouse storage and retrieval optimization via clustering, dynamical systems modeling, and GPU-accelerated routing

@@ 1. Introduction
The system employs clustering on two levels: (1) Order clustering, which reorganizes storage layouts to concentrate frequently co-ordered items and minimize long-term travel distance, and (2) Picking node clustering, which reduces computational complexity in route optimization while maintaining near-optimal results. These techniques are analyzed mathematically in Sections 2 and 3 and evaluated through numerical experiments in Sections 4 and . By continuously adapting storage configurations and applying combinatorial optimization strategies, this approach balances short-term efficiency gains with long-term structural improvements, ensuring that warehouse operations remain effective even amid fluctuating order patterns.
@@
1.1. Related work
A central principle in classical storage location assignment problem (SLAP) is the ABC classification of inventory items according to turnover or demand frequency. In this approach, products are divided into a small number of categories—typically three—such that high-demand (A) items are placed in easily accessible locations, while medium- and low-demand (B and C) items occupy progressively less strategic zones. This strategy has long been recognized as effective for improving picker travel efficiency in picker-to-parts systems [25].
In our study, we adopt $K=3$ clusters to reflect this well-established ABC structure. The intent is not to optimize the number of clusters per se, but to model how a warehouse that follows an ABC-type policy can reorganize its inventory dynamically as demand patterns evolve. The clustering thus serves as an operational analog of ABC zoning under dynamic conditions rather than as a data-driven search for an optimal $K$. For this purpose, we employ the $K$-means algorithm due to its simplicity, interpretability, and computational efficiency. More sophisticated clustering methods (e.g., density-based or hierarchical approaches) could in principle be used, but the aim here is to demonstrate the feasibility and benefits of cluster-driven reorganization rather than to compare clustering techniques. Future work may consider alternative clustering methods within the same framework to explore their potential impact on convergence speed and stability.
Classical research on order picking, as reviewed by de Koster et al. [12], highlights the importance of labor and travel-time considerations in warehouse performance. In contrast, the present work abstracts from such operational constraints—including labor assignment, congestion, and time scheduling—to focus on the structural dynamics of storage-location assignment and product reorganization. Our model assumes unrestricted relocation and omits explicit labor capacity limits, allowing us to isolate and analyze the structural effects of clustering on order consolidation and route optimization. Extending the framework to incorporate labor constraints or relocation delays would represent a natural direction for future research.
While classical studies of picker-to-parts systems emphasize operational batching—where multiple customer orders are grouped to minimize picker travel distance [12]—our approach performs a structural analog of this process. By clustering orders and their associated article types based on spatial and demand similarity, the model effectively performs a dynamic, data-driven form of batching at the storage-location level. This allows reorganization of stock positions that facilitates future co-picking, extending the traditional batching concept from route planning to layout adaptation.
Recent research has also highlighted a growing interest in adaptive storage assignment policies. Rizqi and Chou [27] integrate a mathematical optimization model with discrete-event simulation, using the former to propose storage and picking strategies and the latter to evaluate performance under realistic warehouse dynamics. Other approaches have focused on reinforcement learning: Teck et al. [33] employ deep reinforcement learning (proximal policy optimization, PPO) to optimize rack storage and routing in parts-to-picker systems, while Yerlikaya and Arıkan [35] introduce a multi-criteria class-based storage framework that adapts dynamically to order profiles. In contrast to these simulation- and RL-based frameworks, our contribution provides a tractable mathematical model for picker-to-parts systems that incorporates dynamic product reorganization through clustering. Performance is validated through large-scale numerical experiments supported by GPU-based computation, demonstrating the scalability and stability of the approach.
From a computational perspective, our GPU implementation of the routing component uses the Bellman–Ford algorithm for single-source shortest-path computation (SSSP). While alternative GPU-oriented algorithms such as $\Delta$-stepping can achieve higher asymptotic performance through fine-grained load balancing and reduced synchronization overhead [7], Bellman–Ford offers superior simplicity and generality. Its straightforward parallelization in CUDA enables transparent and reproducible implementation without reliance on external graph-processing libraries. Since routing constitutes only one step within the integrated clustering–RDS framework, Bellman–Ford provides an appropriate and stable baseline, with more specialized GPU SSSP methods left for future large-scale extensions.
Finally, the broader evolution of warehouse research toward data-driven and cyber-physical systems has been comprehensively reviewed by Boysen et al. [6]. They highlight how digitalization, sensor data, and real-time analytics enable adaptive storage assignment and control strategies, often under the umbrella of “Warehousing 4.0.” Our framework aligns with this development by integrating clustering-based reorganization and GPU-accelerated routing into a mathematically transparent model, bridging traditional optimization formulations with scalable, data-oriented computation.
@@ 2.1. The state space
We consider a situation where $N$ types of articles $a_{1},a_{2},\ldots,a_{N}$, are stored in a warehouse at nodes arranged in three-dimensional space representing racks with shelves at different heights. Each shelf on a rack constitutes its own node, and each node is identified with an element in a three-dimensional array having integer coordinates
@@
where $\mathbb{N}=\{0,1,2,\ldots\}$ is the set of natural number, and $n_{x}\times n_{y}\times n_{z}$ is the total number of nodes in the warehouse. We assume there is at most one article type $a_{n}$ at each node (some nodes being empty), and identify each article type $a_{n}$ with the corresponding ordinal $n\in[1,N]$. Empty nodes are assigned article type 0.
For each article type $n\in[1,N]$, let $M_{n}$ be the maximum number of parcels that can be stored at an individual node. We assume this number is independent of which node is being considered; it only depends on what type of article is stored at that node. Introduce the space $\mathcal{A}=[0,N]^{n_{x}\times n_{y}\times n_{z}}$ of three-dimensional arrays. Each array $A=(a_{ijk})_{ijk}\in\mathcal{A}$ corresponds to one particular choice of how to distribute article types in the warehouse. (The array with only zeros represents an empty warehouse with no article types, and the array with only ones is a warehouse that only contains articles of type $a_{1}$.) Each array element $a_{ijk}$ is thus an integer $n$ between 0 and $N$, representing the fact that the node at $(i,j,k)$ contains article type $a_{n}$.
@@
Remark.
In the description above we assume that nodes have integer coordinates and only allow one article type per node. This is a deliberate simplification made for clarity of presentation and ease of interpretation. At the cost of additional bookkeeping, the model can be generalized in two natural ways: to allow nodes with real-valued coordinates, and to allow multiple article types per node (commonly referred to as stock keeping units, or SKUs, in the warehouse literature).
First, multiple SKUs may be assigned to the same node by introducing an additional shelf index $l$, so that
$(i,j,k,l)\in[1,n_{x}]\times[1,n_{y}]\times[1,n_{z}]\times[1,n_{w}]\subset\mathbb{N}^{4},$
where $n_{w}$ denotes shelf capacity. Allocation then corresponds to a standard bin-packing problem, or to a rectangle-packing problem if shelves are treated as two-dimensional footprints [9, 34, 19, 26].
Second, node positions may be generalized to real coordinates by introducing a mapping
$f_{\mathrm{coord}}:\mathbb{N}^{3}\to\mathbb{R}^{3},\qquad(i,j,k)\mapsto\bigl(x(i,j,k),y(i,j,k),z(i,j,k)\bigr),$
so that route optimization is performed on the resulting continuous layout rather than the integer lattice in (1).
While these extensions would make the model more realistic, they are beyond the scope of this paper. Our focus here is on the clustering of orders and picking nodes, and the simplified setting suffices for this purpose. More general formulations, such as multi-SKU nodes or continuous coordinates, are natural directions for future work.
@@ 2.2. Orders
where $m_{n}\in\mathbb{N}$ is the number of parcels of article $a_{n}$ contained in the order. (Here we shall by an order mean any order—individual customer orders or a collection of customer orders—that has been decided should be picked from the warehouse.) Given a state $x=(A,M)$ and an order $o$, the Order Processing algorithm (defined in Algorithm 1 in A below) and a route optimization algorithm (which involves the proposed clustering step based on similarity and proximity, see §2.4) provide the nodes from which the order should be collected, along with the picking route between these nodes. The first part is described by a map $w$ given by
@@
The Check Stock algorithm (Algorithm 3) verifies if there is sufficient stock to fill the order without picking the last parcel on a shelf, i.e., whether $m_{ijk}>w_{ijk}$ for all $(i,j,k)$. If this is not the case, i.e., if $m_{ijk}\leq w_{ijk}$ for some $(i,j,k)$ such that $w_{ijk}\geq 1$, then the Check Stock and Move Articles algorithm of the warehouse optimization routine (Algorithm 4) computes a new warehouse location $(i^{\prime},j^{\prime},k^{\prime})$ at which to restock article $a_{ijk}$ (collecting from the new location after emptying the balance at the old location), and the article array $A$ is updated to $A^{\prime}$ in which $a_{ijk}$ is replaced by 0, and $a_{i^{\prime}j^{\prime}k^{\prime}}$ is changed from 0 to $a_{ijk}$. This gives an updated state
@@ 2.3. Random dynamical systems
To capture the stochastic evolution of warehouse states under repeated picking and resorting, we formulate the system as a random dynamical system (RDS), which provides a compact mathematical framework for modeling how random demand patterns drive successive reorganizations of stored articles over time. Standard mathematical references are Arnold [3] and Strogatz [32].
We use discrete time $n\in\mathbb{N}$, where $\mathbb{N}=\{0,1,2,\ldots\}$ as before is the set of natural numbers. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We take $\Omega$ to be the shift space consisting of sequences $\omega=(o_{1},o_{2},\ldots)$ of orders $o_{n}$, and let $\theta:\Omega\to\Omega$ to be the shift operator $\theta\omega=(o_{2},o_{3},\ldots)$. This models a sequence of orders $\omega\in\Omega$, placed with probability $\mathbb{P}$, coming in to the warehouse to be filled one after the other, with $\theta$ driving the flow of orders. Let $X$ be the state space describing different states of the warehouse. Let $F:\omega\mapsto F_{\omega}$ be a random variable from $\Omega$ to the set of maps on $X$, i.e., for all $\omega\in\Omega$, $F_{\omega}$ is a map $F_{\omega}:X\to X$. Consider the skew product defined by $\Theta(\omega,x)=(\theta\omega,F_{\omega}(x))$. Iterating $\Theta$ gives
@@ 2.4. Routes and clusters
At stopping nodes with more than one picking node, we can pre-determine the order in which picking is carried out—e.g., top to bottom—so this does not have to not included in the route optimization. We will therefore not distinguish between picking routes and routes between stopping nodes.
Figure 1. Each of the top panels show one of three clusters resulting from $K$-means clustering applied to an order with 200 article types given a randomized initial warehouse state $x_{0}$. The nodes in the warehouse have coordinates $(i,j,k)$ with $i,j,k=1,\ldots,10$. Stopping nodes are indicated with color coding describing how many picking nodes a given stopping node contains. Each of the bottom panels show one of three clusters resulting from $K$-means clustering applied to the same order being picked for the 100:th time, that is, given the warehouse state $x_{99}$ as in (5), obtained after $9$ picking iterations. In the top panels, all three clusters are spread out over the entire warehouse with much overlap between clusters. In contrast, the clusters in the bottom panels are concentrated and much better separated.
@@
We remark that we define clusters mathematically as a subset of the picking nodes for a given order, as opposed to dividing the picking nodes from a collection of orders into different groups or clusters. But since an order here can refer to the total collection of individual customer orders to be picked at a certain time our definition covers also the second scenario.
@@
Remark.
In the formulation above, relocations are assumed to be unrestricted, allowing any article to move freely between storage positions at each iteration. This simplifying assumption isolates the structural effects of clustering and routing from operational factors such as labour capacity or relocation delays, which are discussed as possible extensions in Section 6.
@@ 3. Route Optimization
Suppose first that an undirected route with two boundary nodes has been found for each cluster. We determine the computational complexity of finding the best global route going through all clusters along these routes using a combinatorial argument. First, a starting cluster is chosen which can be done in $m$ ways. From that cluster, a global starting point should be determined from the pair of that cluster’s boundary nodes, which can be done in $2$ ways. Next, a second cluster is chosen which can be done in m1m-1 ways. From that cluster, a starting point connecting with the first cluster’s end point should be determined from the second cluster’s pair of boundary nodes, which can be done in $2$ ways. Continuing this way we see that there are $m!\times 2^{m}$ ways to choose a global route with a start and end. Hence, there are half as many global undirected routes, namely $m!\times 2^{m-1}$.
@@
Figure 2. An order with 15 stopping positions clustered into 3 regions (left), and the reduced routes for each cluster represented by two boundary nodes (right). For ease of presentation, stopping nodes are labeled from 0 to $9$, where node 0 corresponds to $(i,j)=(1,1)$ in Figure 1.
@@
Of course, we have been comparing the results of the theorem above to a brute force approach for finding optimal routes without employing sophisticated algorithms to reduce computational complexity. However, using for example the Held & Karp algorithm [21] reduces the time complexity from $\mathcal{O}(n!)$ to $\mathcal{O}(n^{2}\cdot 2^{n})$, which, with $n=12$, is still 3 orders of magnitude higher than the examples above—and this without using a corresponding algorithm to find optimal routes within clusters. Hence, the gains in time complexity from the approximate method are still substantial.
@@ 4. Numerical experiments
    The number of article types is $N=890$, and we assume that the maximum number of articles at a node is $10$ for all article types, so $M_{a_{ijk}}=10$ for all $(i,j,k)$. Thus, the state space $X$ is
@@
    The initial warehouse state $x_{0}$ is randomized with the following constraints: It contains $11$ xyxy-coordinate pairs $(i_{0},j_{0}),(i_{1},j_{1}),\ldots,(i_{10},j_{10})$ such that the nodes $(i_{0},j_{0},k),\ldots(i_{10},j_{10},k)$ are empty for all $k\in[1,10]$. This corresponds to $11$ empty racks, for a total of $110$ empty nodes. A unique article type $a_{ijk}\in[1,890]$ is assigned to each of the remaining $890$ nodes, and we assume that the initial warehouse state is fully stocked, so the balance at each nonempty node $10$.
@@
    An order $o$ to be picked is a collection of $20$ individual customer or purchase orders. A purchase order consists of $10$ article types with between $1$ and $10$ parcels of each type. The $20$ purchase orders are mutually disjoint, so $o$ consists of $200$ article types. In each picking iteration, the $20$ purchase orders are divided into three groups using $K$-means clustering based on where the articles are stored in the warehouse at that time (with $K=3$ reflecting the classical ABC demand classification scheme, as discussed in Section 1.1). Each such group is a cluster. An example can be seen in Figure 1.
@@ 4.1. Experiment 1
Refer to captionRefer to captionRefer to caption
Figure 3. Three snapshots of cluster distributions from experiment 1, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density function plots. The snapshots are taken at time $n$ (i.e., in the language of §2.4 they show the clusters computed for the $n$:th order), with $n=1$ (left), $n=10$ (middle), $n=100$ (right). The orders are the same in every iteration. The figure clearly shows that as time progresses, the separation between cluster centers increases, while the bumps become more articulated due to a decreased variance. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_unperturbed/0_3r1qr1w4.
@@ 4.2. Experiment 2
Refer to captionRefer to captionRefer to caption
Figure 4. Three snapshots of cluster distributions from experiment 2, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density function plots. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $10\%$ on average. Although not as pronounced, we still clearly see an increased separation and decreased variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_static_perturbation/0_t9n5lbme